# Interactive, and Editable Simulation for Autonomous Driving

### v1

@@ III-B Scenario Reconfiguration
The proposed scenario reconfiguration pipeline (refer Fig. 2(c)) can leverage a hierarchical architecture comprising LLM agent(s) as well as rule-based agent(s). The level-1 LLM agent (a.k.a. “Sr. Manager”) parses natural language inputs prompted by the human user(s) to determine the scenario design requirements. These requirements are then passed on to level-2 LLM agents (a.k.a. “Jr. Managers”) to determine the low-level tasks. These specialized tasks are then assigned to the appropriate rule-based agents (a.k.a. “Workers”) to modify the existing scene by searching, adding/removing/updating, or positioning/arranging the requested assets or incorporating specific properties/behaviors to them. This way, the proposed method fosters the creativity of generative AI, while enforcing rule-based bounds to ensure the generation of diverse yet pragmatic scenarios. It is to be noted that the agent hierarchy can be flexibly adjusted depending on the expected complexity and granularity of scenario reconfiguration.
Fig. 1 depicts the high-level aspects of the pipeline. The user inputs a natural language prompt within the designated text field of the graphical user interface (GUI). This prompt is then passed through a prompt engineering function, which adds relevant context and response format to the raw prompt. This ensures robust prompt handling and consistent response formatting for future stages of the hierarchy. As highlighted above, the prompt turns into requirements (processed by level-1 LLM agent), which then turn into specialized tasks (processed by level-2 LLM agents). These specific tasks are executed by rule-based agents to realize the requested scenario.
It is worth mentioning that scenario reconfiguration is an online real-time process, meaning that the simulator continues its functions of physics time-stepping, rendering, data transfer via application programming interfaces (APIs), accepting inputs and providing feedback via the graphical user interface (GUI) or the human-machine interface (HMI), etc. Additionally, being an online process, the pipeline supports multi-stage commands while preserving context from the previous prompts. This is especially useful when sequentially editing the scenario while accepting or rejecting previous changes.
@@ IV-A Scene Reconstruction
The following metrics were used to assess the quality and performance of the scene/asset reconstructions:
@@
PSNR: Peak signal-to-noise ratio (PSNR) quantifies the reconstruction fidelity by comparing the pixel-wise similarity between the rendered and the ground truth images. It focuses on pixel intensity differences and is derived from the mean squared error (MSE) or $L_{2}$ distance between the two images.
$\mathrm{PSNR}(x,y)=10\cdot log_{10}\left(\frac{\mathrm{MAX}_{I}^{2}}{\mathrm{MSE}(x,y)}\right)$ (4)
where, $\mathrm{MAX}_{I}$ is the maximum possible pixel value (e.g., $2^{\mathrm{n}}$ for $n$-bit images). A higher PSNR indicates better reconstruction quality (lower error). PSNR can measure even small pixel-level differences but may fail to capture perceptual distortions such as blurriness or loss of structural information.
TABLE I: Scene Reconstruction Benchmarking
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
$\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$
3DGS 23.71 2.07 0.78 0.09 0.20 0.09 63.55 23.84
NeRF 18.35 2.00 0.41 0.16 0.56 0.16 0.57 0.27
TABLE II: Scene Reconstruction Performance Evaluation
CU-ICAR 1420 23.71 0.78 0.20 63.55
CGEC 1241 23.47 0.74 0.32 61.96
AuE Lab 545 35.10 0.97 0.17 69.36
Footpath (Lit) 70 32.20 0.93 0.09 66.09
Footpath (Unlit) 263 31.82 0.90 0.22 68.27
Passenger Car 971 24.93 0.88 0.19 79.19
Traffic Cone 530 35.11 0.95 0.06 62.65
Pedestrian Sign 1102 31.41 0.94 0.08 64.40
SSIM: Structural similarity index measure (SSIM) evaluates the perceptual structural similarity between the rendered image and the ground truth by comparing luminance $l(x,y)=\frac{2\mu_{x}\mu_{y}+c_{1}}{\mu_{x}^{2}+\mu_{y}^{2}+c_{1}}$, contrast $c(x,y)=\frac{2\sigma_{x}\sigma_{y}+c_{2}}{\sigma_{x}^{2}+\sigma_{y}^{2}+c_{2}}$, and structure $s(x,y)=\frac{\sigma_{xy}+0.5c_{2}}{\sigma_{x}\sigma_{y}+0.5c_{2}}$. It is designed to align with human visual perception.
$\mathrm{SSIM}(x,y)=[l(x,y)]^{\alpha}\cdot[c(x,y)]^{\beta}\cdot[s(x,y)]^{\gamma}$ (5)
@@
where, $\{\alpha,\beta,\gamma\}>0$ weigh the relative importance of $l$, $c$, and $s$ respectively. SSIM ranges from -1 to 1, indicating the degree of correlation between the pair of images (1 means perfect similarity). It captures perceptual quality better than PSNR by focusing on local image structures, which are critical for evaluating how well a method reconstructs edges, textures, and scene consistency.
@@
LPIPS: Learned perceptual image patch similarity (LPIPS) is a deep-learning-based perceptual metric that compares high-level features extracted from a neural network (e.g., AlexNet in our case) rather than pixel-wise differences. It measures how perceptually similar two images are by computing the $L_{2}$ distance between their feature activations at multiple layers of a pre-trained neural network.
$\mathrm{LPIPS}(x,y)=\sum_{l}\frac{1}{H_{l}\cdot W_{l}}\sum_{h,w}\\ \left\|w_{l}\odot\left(\phi^{l}(x)_{h,w}-\phi^{l}(y)_{h,w}\right)\right\|_{2}^{2}$ (6)
LPIPS score ranges from 0 (perfect similarity) to 1 (completely different). It evaluates perceptual quality at a high level, capturing differences in texture, details, and global image structure that are not detected by PSNR or SSIM.
@@
FPS: Frames per second (FPS) is a performance measure to assess the rendering rate [^3] All experiments were conducted using a single laptop PC with 12th Gen Intel Core i9-12900H 2.50 GHz CPU, NVIDIA GeForce RTX 3080 Ti GPU, and 32.0 GB RAM.. It measures the number of frames $\mathrm{n_{frames}}$ rendered within a predefined time interval $T$.
$\mathrm{FPS}=\frac{\mathrm{n_{frames}}}{T}$ (7)
FPS values can range from 0 (no update) to upwards of 30 (real-time), with $>$60 FPS being a good benchmark for autonomy-oriented simulations.
@@
We first benchmark two different photorealistic reconstruction methods, viz. NeRF and 3DGS. Table I presents the reconstruction results for the CU-ICAR scene (used as a benchmark). 3DGS outperforms NeRF by a large margin: not only are the quality metrics for 3DGS better than NeRF (PSNR +29.21%, SSIM +90.24%, LPIPS -64.29%), but 3DGS is over 100 times faster than NeRF for the same scene. The choice of 3DGS is thus justified, and further results will assume 3DGS reconstruction, unless mentioned otherwise.
Following the benchmarking experiment, we analyzed the performance of 3D reconstruction across various scenes and assets. Here, we considered scene variations in terms of their type (indoor/outdoor), scale (small/large), and lighting (lit/unlit). Diversity in assets was considered in terms of their materials (color, texture, opacity, albedo, etc.) and scale (small/large). Table II presents the data requirements, reconstruction quality, and rendering rate for 5 scenes and 3 assets.
It was observed that large outdoor scenes typically required upwards of 1k frames for effective reconstruction. Smaller (indoor) scenes, on the other hand, could be reconstructed with just half the data, but their depth estimation was observed to be slightly inaccurate, generating a few stray artifacts. This issue (inaccurate depth mapping leading to stray artifacts) was more prominently observed for the footpath scenes, where the data quantity was severely reduced. We also noticed that a well-lit footpath scene required significantly ($\sim$4$\times$) less data than a similar footpath scene, which was poorly lit.
It was inferred that the scale or type of scene/asset did not directly govern the reconstruction fidelity. Instead, the quality (coverage, lighting, resolution, frame-rate, etc.) of the data mattered the most. However, it is also arguable that collecting good-quality data for smaller assets and indoor scenes is much easier than their larger and outdoor counterparts.
IV-B Scenario Reconfiguration
@@
Figure 4: LLM-guided reconfiguration: Digital twin of the CU-ICAR campus being reconfigured to (a) daytime; (b) nighttime; (c) foggy weather; (d) rainfall; (e) snowfall; (f) clear weather; (g-i) sequentially add maintenance work assets; (j) include a parked car; (k) add a 1:5 scale mobile robot controlled to perform a specific maneuver; (l) include a jaywalking pedestrian.
The design of experiments to analyze the LLM-guided scenario reconfiguration pipeline was split into benchmarking (refer Table III) and evaluation (refer Table IV) categories.
In terms of benchmarking, we employed 3 different open-access LLMs, viz. Mistral 0.2, Llama 3.1, and Gemma 2. The choice of these models was primarily driven by their open accessibility, though it is worth mentioning that the proposed framework is capable of using any other models or APIs (both open as well as commercial). We consider the generalizability and repeatability of the model as our primary metrics for benchmarking, in addition to the model’s disk size occupancy, number of parameters, and inference time (from prompt input to scenario update). Here, repeatability is defined as the ability to produce the same result/response for 100 trials across 7 tasks with a direct prompt. Whereas, generalizability is defined as the ability to reconfigure the scenario satisfactorily for 100 trials across 7 tasks with 4 grades of prompts.
The prompt types ranged from explicit/direct to incorrect/erroneous, with a total of 4 different gradations as defined below:
@@
  • Indirect Prompt: These prompts clearly express the user’s intent but do not provide exact asset names, numeric values, or similar information. Such commands can be expected from users with a fair amount of knowledge, but who are not necessarily aware of specific terminology or conventions.
  • Vague Prompt: These prompts express the user’s intent minimally, and some things are left for interpretation by the model. Such commands can be expected from users with superficial or no knowledge.
@@
TABLE III: Scenario Reconfiguration Benchmarking
LLM Size (GB) Params. (#) Gen. (%) Rep. (%) Time (s)
@@
TABLE IV: Scenario Reconfiguration Performance Evaluation
@@
The tasks included (i) searching for specific assets or properties, (ii) addition of assets, (iii) removal of assets, (iv) positioning the assets according to spatial cues, (v) moving the assets, (vi) arranging homogeneous or heterogeneous assets, and (vii) modifying appearance of the scene.
As expected, models with more parameters occupied higher disk space and incurred larger inference times. Particularly, it was observed that for every 1B additional parameters, the model size increased by $\sim$0.5 GB, and its inference time by $\sim$1 second. However, for the objective of scenario reconfiguration, each increment in the parameter space improved the model performance (generalizability and repeatability) significantly. This performance improvement outweighed the slight increase in model size, given modern digital storage systems. Similarly, a slightly higher inference time (order of milliseconds) could not be compared to a human user hand-crafting the scenario (order of hours or days).
@@
In terms of model-specific performance, it was observed that Mistral 0.2 provided highly overconfident responses, resulting in high repeatability without much generalizability. Llama 3.1 and Gemma 2 performed comparably, but the latter was slightly better and generated more elegant responses. Consequently, further results will assume the Gemma 2 model, unless mentioned otherwise.
@@
A detailed performance evaluation of the Gemma 2 model indicated that direct prompts resulted in over 90% satisfaction for the majority of the tasks, except for arrangement, which was an inherently complex task. Reducing the quality of prompts to indirect structuring resulted in degradation of positioning and moving tasks, without significantly affecting others. Further degrading the prompt quality to vague structuring resulted in a significant degradation of satisfaction across all the tasks, with search and appearance tasks being affected the most. Finally, adulterating the prompts with grammatical and/or structural errors resulted in a noticeable degradation of satisfaction across all the tasks, with arrangement and search tasks being affected the most.
@@
We also conducted experiments with varying complexities of prompts and analyzed them qualitatively. These started with simple tasks like changing the time of day (refer Fig. 4(a-b)) or weather conditions (refer Fig. 4(c-f)). Next, we performed multi-stage scenario editing using sequential prompting (refer Fig. 4(g-i)) without repeating earlier information, to assess the context-retention ability. Finally, we also prompted complex commands involving several types of tasks to be executed simultaneously (refer Fig. 4(j-l)) to assess the task management ability. It was interesting to see the spatial awareness (e.g., implicit parking location in Fig. 4(j)) of the model, as well as its context-aware creativity (e.g., adding a pedestrian sign without being explicitly asked in Fig. 4(l)).

### v2

@@ III-B Scenario Reconfiguration
The proposed scenario reconfiguration pipeline (refer Fig. 2(c)) can leverage a hierarchical architecture comprising LLM agent(s) as well as rule-based agent(s). The level-1 LLM agent (a.k.a. “Sr. Manager”) parses natural language inputs prompted by the human user(s) to determine the scenario design requirements based on specific keywords. An engineered prompt adds relevant context (based on the task/scene description, localization, and the asset library) as well as response format and error-handling instructions to the raw prompt. This ensures robust prompt handling and consistent response formatting for further levels of the hierarchy. The requirements are then passed on to level-2 LLM agents (a.k.a. “Jr. Managers”) to determine the low-level tasks. Again, an engineered prompt adds relevant context, response format, and error-handling instructions to the raw query. These specialized tasks are then assigned to the appropriate rule-based agents (a.k.a. “Workers”) to modify the existing scene by searching for, adding/removing/updating, or positioning/arranging the requested assets, or by incorporating specific properties/behaviors into them. The rule-based agents also implement pre-execution validation and error-handling mechanisms to avoid ill-posed scenario reconfiguration. This way, the proposed method fosters the creativity of generative AI, while enforcing rule-based constraints to ensure the generation of diverse yet pragmatic scenarios. It is worth noting that the agent hierarchy can be flexibly adjusted based on the expected complexity and granularity of scenario reconfiguration. Fig. 5 presents a simple and intuitive worked example of scenario reconfiguration.
Figure 5: Worked example of scenario reconfiguration: (a) User prompt “make it rain” being handled by level-1 LLM agent, (b) scenario design requirements being parsed by level-2 LLM agent, and (c) specialized tasks being implemented by rule-based agent to modify the existing scene.
It is worth noting that scenario reconfiguration is an online, real-time process, meaning that the simulator continues to perform physics time-stepping, rendering, data transfer via application programming interfaces (APIs), accepting inputs, and providing feedback via the graphical user interface (GUI) or the human-machine interface (HMI), etc. Additionally, being an online process, the pipeline supports multi-stage commands while preserving context from the previous prompts. This is especially useful when sequentially editing the scenario while accepting or rejecting previous changes.
The proposed optional VLM-based visual enhancement pipeline leverages conditional latent diffusion models (e.g., SDXL). Particularly, the prompt engineering ensures that shading, lighting, reflections, shadows, and volumetric effects are blended throughout the scene without affecting the pose, scale, or other attributes of the scene/assets. It is also worth noting that this optional image enhancement can be performed online (e.g., to modify the rasterized output from the simulator, albeit at the cost of reduced frame rate) or offline (e.g., to batch-process synthetic datasets).
@@ IV-A Scene Reconstruction
The following metrics [19, 20] were used to assess the quality and performance of the scene/asset reconstructions:
CCD/CMD: Cloud-cloud distance (CCD) and cloud-mesh distance (CMD) quantify geometric reconstruction accuracy by measuring the point-wise (CCD) or point-to-surface (CMD) nearest-neighbor Euclidean distances (in meters) between a reconstructed point cloud and a reference point cloud (CCD), or reference mesh (CMD). Lower CCD/CMD values indicate better geometric alignment. CCD and CMD are sensitive to local geometric errors and can capture fine-scale surface discrepancies; however, they may be influenced by point density, noise, and alignment quality, and do not directly reflect the perceptual or semantic correctness.
@@
PSNR: Peak signal-to-noise ratio (PSNR) quantifies the reconstruction fidelity by comparing the pixel-wise similarity between the rendered and the ground truth images. It focuses on pixel intensity differences and is derived from $L_{2}$ distance between the two images. Higher PSNR indicates better reconstruction quality. PSNR can measure even small pixel-level differences but may fail to capture perceptual distortions such as blurriness or loss of structural information.
@@
SSIM: Structural similarity index measure (SSIM) evaluates the perceptual structural similarity between the rendered image and the ground truth by comparing luminance, contrast, and structure. It is designed to align with human visual perception. SSIM ranges from -1 to 1, indicating the degree of correlation between the pair of images (1 corresponds to perfect similarity). It captures perceptual quality better than PSNR by focusing on local image structures, which are critical for evaluating how well a method reconstructs edges, textures, and scene consistency.
@@
LPIPS: Learned perceptual image patch similarity (LPIPS) is a deep-learning-based perceptual metric that compares high-level features extracted from a neural network (e.g., AlexNet in our case) rather than pixel-wise differences. It measures perceptual similarity between two images by computing the $L_{2}$ distance between their feature activations across multiple layers of a pre-trained neural network. LPIPS score ranges from 0 (perfect similarity) to 1 (completely different). It evaluates perceptual quality at a high level, capturing differences in texture, details, and global image structure.
@@
FPS: Frames per second (FPS) is a performance measure to assess the rendering rate [^3] All experiments were conducted using a single laptop PC with 12th Gen Intel Core i9-12900H 2.50 GHz CPU, NVIDIA GeForce RTX 3080 Ti GPU, and 32.0 GB RAM.. It measures the number of frames rendered within a predefined time interval. FPS values can range from 0 (no update) to upwards of 30 (real-time), with >>60 FPS being a good benchmark for autonomy-oriented simulations.
@@
We first benchmark two different photorealistic reconstruction methods, viz. NeRF and 3DGS. Table II presents the reconstruction results for the CU-ICAR scene (used as a benchmark). 3DGS outperforms NeRF by a large margin: not only are the quality metrics for 3DGS better than NeRF (PSNR +29.21%, SSIM +90.24%, LPIPS -64.29%), but 3DGS is over 100 times faster than NeRF for the same scene. The choice of 3DGS is thus justified, and further results will assume 3DGS reconstruction, unless mentioned otherwise. Furthermore, as discussed earlier, we compare and propose to couple 3DGS-based photorealism (higher PSNR/SSIM and lower LPIPS) with PSR-based geometrical precision (lower CCD/CMD) and dynamic interactability (continuous surface representation) to complement each other.
@@
Figure 6: LLM-guided reconfiguration: Digital twin of the CU-ICAR campus being reconfigured to (a) daytime; (b) nighttime; (c) foggy weather; (d) rainfall; (e) snowfall; (f) clear weather; (g-i) sequentially add maintenance work assets; (j) add a parked car at the specified location; (k) add a mobile robot controlled to perform a specific maneuver; (l) include a jaywalking pedestrian.
Following the benchmarking experiment, we analyzed the performance of 3D reconstruction. Here, we considered scene variations in terms of their setting (indoor/outdoor), scale (small/large), and lighting (lit/unlit). Diversity in assets was considered in terms of their color, texture, opacity, albedo, and scale. Table III presents the data requirement, reconstruction quality, and FPS for 5 scenes and 3 assets depicted in Fig. 4.
[…]
It was inferred that the scale or type of scene/asset did not directly govern the reconstruction fidelity. Instead, the quality (coverage, lighting, resolution, frame rate, etc.) of the data mattered most. However, it is also arguable that collecting good-quality data for smaller assets and indoor scenes is much easier than their larger and outdoor counterparts.
TABLE II: Scene Reconstruction Benchmarking
Method CCD/CMD $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
NeRF [4] 1.83e-1 18.35 0.41 0.56 0.57
3DGS [5] 6.54e-2 23.71 0.78 0.20 63.55
PSR [2] 3.76e-4 8.74 0.32 0.89 144.26
3DGS+PSR (Ours) 3.76e-4 23.71 0.78 0.20 61.72
TABLE III: Scene Reconstruction Performance Evaluation
CU-ICAR 1420 23.71 0.78 0.20 63.55
CGEC 1241 23.47 0.74 0.32 61.96
AuE Lab 545 35.10 0.97 0.17 69.36
Lit Footpath 70 32.20 0.93 0.09 66.09
Unlit Footpath 263 31.82 0.90 0.22 68.27
Passenger Car 971 24.93 0.88 0.19 79.19
Traffic Cone 530 35.11 0.95 0.06 62.65
Pedestrian Sign 1102 31.41 0.94 0.08 64.40
IV-B Scenario Reconfiguration
The design of experiments to analyze the LLM-guided scenario reconfiguration pipeline was split into benchmarking (refer Table IV) and evaluation (refer Table V) categories.
In terms of benchmarking, we employed 3 different open-access LLMs, viz. Mistral 0.2, Llama 3.1, and Gemma 2. The choice of these models was primarily driven by their open accessibility, though it is worth mentioning that the proposed framework is capable of using any other models or APIs (both open as well as commercial). We consider the generalizability and repeatability of the model as our primary metrics for benchmarking, in addition to the model’s disk space occupancy, number of parameters, and inference time (from prompt input to scenario update). Here, repeatability is defined as the ability to produce the same result/response for 100 trials across 7 tasks with a direct prompt. Whereas, generalizability is defined as the ability to reconfigure the scenario satisfactorily for 100 trials across 7 tasks with 4 grades of prompts.
The prompt types ranged across 4 different gradations:
@@ IV-B Scenario Reconfiguration
  • Indirect Prompt: These prompts clearly express the user’s intent but do not provide exact asset names, numeric values, or other information. Such commands can be expected from users with a fair amount of knowledge, but who are not necessarily aware of specific terminology or conventions.
  • Vague Prompt: These prompts express the user’s intent minimally, and some things are left for interpretation by the LLM. Such commands can be expected from users with superficial or no knowledge.
@@
TABLE IV: Scenario Reconfiguration Benchmarking
LLM Size (GB) Params. (#) Gen. (%) $\uparrow$ Rep. (%) $\uparrow$ Time (s) $\downarrow$
@@
TABLE V: Scenario Reconfiguration Performance Evaluation
@@
The tasks included (i) searching for specific assets or properties, (ii) addition of assets, (iii) removal of assets, (iv) positioning the assets according to spatial cues, (v) moving the assets, (vi) arranging homogeneous or heterogeneous assets, and (vii) modifying appearance of the scene/assets.
The reconfiguration capabilities of various LLMs were assessed based on whether they satisfactorily accomplished the intended task (a single task at a time). These assessments were based on the updated simulation states after scenario reconfiguration, which were manually analyzed against the intended task stated in the prompt (incorrect/inaccurate execution of the task or inability to comprehend the prompt were both considered failures).
@@
In terms of model-specific performance (refer Table IV), it was observed that Mistral 0.2 provided highly overconfident responses, resulting in high repeatability without much generalizability. Llama 3.1 and Gemma 2 performed comparably, but the latter was slightly better and generated more elegant responses. Consequently, subsequent results will assume the Gemma 2 model, unless mentioned otherwise.
@@
A detailed performance evaluation of the Gemma 2 model (refer Table V) indicated that direct prompts resulted in over 90% satisfaction for the majority of the tasks, except for arrangement, which was an inherently complex task. Reducing the quality of prompts to indirect structuring resulted in degradation of positioning and moving tasks, without significantly affecting others. Further degrading the prompt quality to vague structuring resulted in a significant degradation across all the tasks, with search and appearance tasks being affected the most. Finally, adulterating the prompts with errors resulted in a noticeable degradation across all the tasks, with arrangement and search tasks being affected the most.
(a) Day: Baseline
(b) Day: Enhanced
(c) Night: Baseline
(d) Night: Enhanced
(e) Fog: Baseline
(f) Fog: Enhanced
(g) Rain: Baseline
(h) Rain: Enhanced
(i) Snow: Baseline
(j) Snow: Enhanced
(k) Maintenance: Baseline
(l) Maintenance: Enhanced
Figure 7: VLM-guided enhancement: Digital twin of the CU-ICAR campus being visually enhanced across (a-b) daytime; (c-d) nighttime; (e-f) foggy weather; (g-h) rainfall; (i-j) snowfall; (k-l) maintenance work. Notice appearance enhancement, including shading, lighting, reflections, shadows, and blending of volumetric effects, without affecting the pose or scale of the scene/assets.
@@
We also conducted experiments with varying levels of prompt complexity and qualitatively analyzed them. These started with simple tasks like changing the time of day (refer Fig. 6(a-b)) or weather conditions (refer Fig. 6(c-f)). Next, we performed multi-round scenario editing using sequential prompting (refer Fig. 6(g-i)) without repeating earlier information, to assess the context-retention ability. Finally, we also prompted complex commands involving several types of tasks to be executed simultaneously (refer Fig. 6(j-l)) to assess the task management ability. It was interesting to see the spatial awareness of the model (e.g., implicit parking location in Fig. 6(j)), as well as its context-aware creativity (e.g., adding a pedestrian sign without being explicitly asked in Fig. 6(l)).
In terms of model size, parameters, and inference time, it was observed that for every 1B additional parameters, the model size increased by $\sim$0.5 GB, and its inference time by $\sim$1 second. However, for the objective of scenario reconfiguration, each increment in the parameter space significantly improved model performance (generalizability and repeatability). This performance improvement outweighed the slight increase in model size, given modern digital storage systems. Similarly, slightly higher inference time (order of milliseconds) could not be compared to a manual hand-crafting (order of hours/days).
TABLE VI: Visual Enhancement Benchmarking
Method FID $\downarrow$ KID $\downarrow$ CMMD $\downarrow$ AIDP $\downarrow$ FPS $\uparrow$
Baseline 13.33 0.44 0.57 0.55 63.55
Enhanced 13.25 0.39 0.48 0.30 06.68
Ultimately, we benchmarked the optional VLM-based visual enhancement module. Particularly, we compared the baseline image frames with their enhanced counterparts qualitatively (refer Fig. 7 for some examples) as well as quantitatively (refer Table VI) across the following metrics [20]:
FID: Fréchet Inception distance (FID) evaluates the realism of generated images by comparing the distributions of features extracted from generated and real images using a pretrained Inception network. It measures the distance between two multivariate Gaussian distributions characterized by their means and covariances. Lower FID indicates closer alignment to real image statistics (higher realism). While FID captures high-level semantic and perceptual differences better than pixel-wise metrics, it can be sensitive to sample size and feature extractor bias.
KID: Kernel Inception distance (KID) measures the similarity between generated and real image distributions using a kernel-based maximum mean discrepancy (MMD) computed on Inception features. Unlike FID, KID provides an unbiased estimator for finite sample sizes. Lower KID values indicate better agreement between generated and real image distributions. KID captures semantic discrepancies but, like FID, does not directly assess image fidelity at the pixel level.
CMMD: CLIP-MMD (CMMD) quantifies the alignment between generated and reference image distributions by computing the MMD in a shared CLIP embedding space. By leveraging CLIP’s multi-modal representations, this metric captures semantic and perceptual similarity that aligns well with human judgment. Lower CMMD indicates closer distributional similarity, though results depend on the chosen CLIP model and may overlook fine-grained visual artifacts.
AIDP: AI detector probability (AIDP) estimates the likelihood that an image is generated by an artificial model, typically produced by a trained classifier distinguishing real versus synthetic content. Lower probabilities suggest more realistic images, while higher probabilities indicate detectable generative artifacts. Although helpful in assessing detectability, these scores are model-dependent, can change over time, and are proxy measures of visual quality or semantic correctness.
Results consistently indicated lower AIDP values for the enhanced images, indicating higher visual realism or quality compared to their baseline counterparts. Furthermore, the FID, KID, and CMMD proxy metrics also confirmed this trend, indicating that the enhanced image distribution was closer to the distribution of “real” images. However, it is worth mentioning that due to the added latency of the VLM inference, the simulation FPS dropped by an order of magnitude (from 63.55 to 6.68), though this bottleneck can be addressed relatively easily by leveraging high-performance computing (HPC) resources (e.g., >>65 FPS with L40S GPU).